OpenStudy (anonymous):

Solve 2x2 + 20x = −38.

OpenStudy (luffingsails):

Can you use the quadratic formula?

OpenStudy (anonymous):

Use the quadratic formula, applying it to:\[ \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] And then simplify.

OpenStudy (anonymous):

SIMPLIFY

OpenStudy (bee_see):

write the the problem to ax^2+bx+c=0

OpenStudy (nevermind_justschool):

Wow, it says the question was posted six days ago...